OpenStudy (anonymous):

Classify the system and determine the number of solutions. (Check all that apply.) 7x - y = -11 3y = 21x + 33 independent inconsistent infinitely many solutions consistent no solution dependent

OpenStudy (tkhunny):

Find a Solution? Consistent. Now what?

OpenStudy (anonymous):

its just consistent?

OpenStudy (anonymous):

i get so confused with these & i only have 6 mins left on my test :(

OpenStudy (tkhunny):

I didn't say it was anything about dependent or independent or whether or not it had a solution. The difference between consistent and inconsistent is whether you can find a solution. Do that. See if you can find a solution. If you can, it's consistent. Then, we can worry about dependent or independent.

OpenStudy (anonymous):

theres no solution

OpenStudy (tkhunny):

What test are we cheating on? 1) 7x - y = -11 2) 3y = 21x + 33 Change the top equation a little 1) 7x - y = -11 Add y and add 11 1) y = 7x + 11 Multiply by 3 1) 3y = 21x + 33 Now, copy down equation #2. 2) 3y = 21x + 33 This gives it all away.

OpenStudy (tkhunny):

There are infinitely many solutions. How did you manage to find none of them?